# LeOCLR: Leveraging Original Images for Contrastive Learning of Visual Representations

### v1

@@ 4 Experiments
5 Ablation Studies
In this section, we conduct further analysis of our approach using another contrastive instance discrimination approach SimCLR[7] to explore how our approach will perform within this end-to-end framework. Also, we perform studies on the benchmark datasets STL-10 and CIFAR-10 with a different backbone (Resnet18) to check the consistency of our approach with other datasets and backbones. Furthermore, we employ a random crop test to simulate natural transformations, such as variations in scale or occlusion of objects appearing in the image, in order to conduct further analysis on the robustness of features learned by our approach, LeOCLR. In addition, we compare our approach with vanilla MoCo when manipulating their data augmentation to see which model’s performance is more affected by removing some of the data augmentation. Finally, we use different fine-tuning settings to check which model learns better and faster.
@@ 5 Ablation Studies
We use an end-to-end framework, where the two encoders $f_{q}$ and $f_{k}$ are updated via backpropagation, to train a model with our approach for 200 epochs and 256 batch size. Subsequently, we perform a linear evaluation of our model against SimCLR, which uses an end-to-end mechanism. As shown in Tab. 4, our approach outperforms vanilla SimCLR by a significant margin of 3.5%, demonstrating its suitability for integration with various contrastive learning frameworks.
@@
In Tab. 5, we evaluate our approach on different datasets (STL-10 and CIFAR-10) using another backbone, namely ResNet18, to ensure its consistency across various backbones and datasets. We pre-trained both models (Vanilla MoCo-v2 and LeOCLR) for 800 epochs on both datasets and then conducted a linear evaluation for both models. Our approach demonstrates superior performance on both datasets compared to vanilla MoCo-v2, achieving accuracies of 5.12% and 5.71% on STL-10 and CIFAR-10, respectively.
@@
MoCo-v2 [8] 67.5% 63.2%
CLSA [39] 69.4% -
@@
In Tab. 6, we reported the top-1 accuracy for vanilla MoCo-v2 and our approach after 200 epochs on ImageNet. Tab. 6 shows two testing methods: center crop test similar to [7, 8]: images are resized to 256 pixels along the shorter side using bicubic resampling, after which a $224\times 224$ center crop is applied. The second test is a random crop, where the image is resized to $256\times 256$ but randomly cropped and resized to $224\times 224$. We took the MoCo-v2 center crop result directly from [8], while the random crop result was not reported. Therefore, we replicated the MoCo-v2 with the same hyperparameters used in the original paper to report the center crop, ensuring a fair comparison. According to the results, the performance of MoCo-v2 dropped by 4.3% with random cropping, whereas our approach experienced a smaller drop of 2.8%. This suggests that our approach learns better semantic features, as it demonstrates greater invariance to natural transformations such as occlusion and variations in object scales. Also, we compare the performance of CLSA [39] with our approach because we have the same performance after 800 epochs (see Tab. 1. Note that the CLSA approach uses multi-crop (i.e., five strong and two weak augmentations), while our approach only uses two random crops and the original image. As shown in Tab. 5 LeOCLR outperforms the CLSA approach by 2.3% after 200 epochs on ImageNet-1K.
Figure 4: Decrease in top-1 accuracy (in % points) of LeOCLR and our own reproduction of Vanilla MoCo at 200 epochs, under linear evaluation on ImageNet. $R\_Grayscale$ means to remove the grayscale augmentations, and $R\_color$ removes color jitter with grayscale augmentations.
Contrastive instance discrimination approaches are sensitive to the choice of image augmentations [16]. Thus, we do further analysis of our approach against Moco-v2 [8]. These experiments aim to see which model learns better semantic features and creates robust representation under different data augmentations. As shown in Fig. 4, both models are affected by removing some data augmentations. However, our approach shows a more invariant representation and Less performance impact due to transformation manipulation than vanilla MoCo-v2. For example, when we apply only random cropping augmentation, the performance of vanilla MoCo-v2 is reduced by 28 points (i.e., from 67.5% baseline to 39.5% only random cropping), while our approach reduces only 25 points (i.e., from 71.7% baseline to 46.6% only random cropping). This means our approach learns better semantic features and creates better representation for the given objects than vanilla MoCo-v2.
(a) Top-1 accuracy
(b) Top-5 accuracy
@@
Tab. 2 presented in Sec. 4, we fine-tune the representation over the 1% and 10% ImageNet splits from [7] with ResNet-50 architecture. In the ablation study, we compare the fine-tuned representation of our approach and reproduced vanilla MoCo-v2[8] over 1%, 2%, 5%, 10%, 20%, 50%, and 100% of the ImageNet dataset as in [20, 16]. In this setting, we observed that tuning a LeOCLR representation always outperforms vanilla MoCo-v2. For instance, Fig. 5 (a) shows that LeOCLR fine-tuned with 10% of ImageNet labeled data performed better than Vanilla Moco-v2 [8] fine-tuned with 20% of labeled data. This means that our approach is suitable in case we have small labeled data for downstream task than vanilla MoCo-v2.
@@ 6 Conclusion
In this paper, we introduce a new SSL approach to improve contrastive instance discrimination representation learning. Our approach alleviates discarding semantic features while attracting two views containing distinct semantic content by incorporating the original image in training. Our approach consistently enhances representation learning for contrastive instance discrimination across different benchmark datasets and various mechanisms, such as momentum contrast and end-to-end methods. In linear evaluation, we achieved an accuracy of 76.2% on ImageNet after 800 epochs, outperforming several SOTA SSL approaches.
Through extensive ablation studies and experiments, we have demonstrated the robustness and invariance of our method to different backbones and datasets. These findings suggest that our method could be a promising candidate for adoption in various settings where semi-supervised learning methods are employed, extending beyond the contexts considered in this paper.

### v2

@@ 4 Experiments and Results
Object Detection Task: To further evaluate the transferability of the learned representation, we fine-tune the model on PASCAL VOC object detection. We use similar settings as in MoCo (Chen et al., 2020b), where we finetune on the VOC07+12 trainval dataset using Faster R-CNN with a R50-C4 backbone and evaluate on VOC07 test dataset. We fine-tuned for 24k iterations ($\approx$ 23 epochs). Our method outperforms all compared approaches, as shown in Table 4. This is because it excels at capturing semantic features compared to the baseline (MoCo-v2) and other solutions tackling the same problem, which leads to superior performance on object detection and other related tasks.
Table 4: Results (Average Precision) for PASCAL VOC object detection using Faster R-CNN with ResNet50-C4.
Approach $AP_{50}$ $AP$ $AP_{75}$
MoCo-v2 (Chen et al., 2020b) 82.5% 57.4% 64%
CLSA (Wang & Qi, 2022) 83.2% - -
SCFS (Wang & Qi, 2022) 83% 57.4% 63.6%
LeOCLR(ours) 83.2% 57.5% 64.2%
5 Ablation Study
In this section, we conduct further analysis of our approach using another contrastive instance discrimination approach, SimCLR(Chen et al., 2020a), to explore how our approach will perform within this end-to-end framework. Also, we conduct studies on the benchmark datasets STL-10 and CIFAR-10 with a different backbone (Resnet18) to check the consistency of our approach with other datasets and backbones. Furthermore, we employ a random crop test to simulate natural transformations, such as variations in scale or occlusion of objects appearing in the image, in order to conduct further analysis on the robustness of features learned by our approach, LeOCLR. In addition, we compare our approach with vanilla MoCo when manipulating their data augmentation to see which model’s performance is more affected by removing some of the data augmentation. Finally, we use different fine-tuning settings to check which model learns better and faster.
@@ 5 Ablation Study
We use an end-to-end framework, where the two encoders $f_{q}$ and $f_{k}$ are updated via backpropagation, to train a model with our approach for 200 epochs and 256 batch size. Subsequently, we perform a linear evaluation of our model against SimCLR, which uses an end-to-end mechanism. As shown in Table 5, our approach outperforms vanilla SimCLR by a significant margin of 3.5%, demonstrating its suitability for integration with various contrastive learning frameworks.
@@
In Table 6, we evaluate our approach on different datasets (STL-10 and CIFAR-10) using another backbone, namely ResNet18, to ensure its consistency across various backbones and datasets. We pre-trained both models (Vanilla MoCo-v2 and LeOCLR) for 800 epochs on both datasets and then conducted a linear evaluation for both models. Our approach demonstrates superior performance on both datasets compared to vanilla MoCo-v2, achieving accuracies of 5.12% and 5.71% on STL-10 and CIFAR-10, respectively.
@@
MoCo-v2 (Chen et al., 2020b) 67.5% 63.2%
CLSA (Wang & Qi, 2022) 69.4% -
@@
In Table 7, we reported the top-1 accuracy for vanilla MoCo-v2 and our approach after 200 epochs on ImageNet. Table 7 shows two testing methods: center crop test similar to (Chen et al., 2020a; b): images are resized to 256 pixels along the shorter side using bicubic resampling, after which a $224\times 224$ center crop is applied. The second test is a random crop, where the image is resized to $256\times 256$ but randomly cropped and resized to $224\times 224$. We took the MoCo-v2 center crop result directly from (Chen et al., 2020b), while the random crop result was not reported. Therefore, we replicated the MoCo-v2 with the same hyperparameters used in the original paper to report the center crop, ensuring a fair comparison. According to the results, the performance of MoCo-v2 dropped by 4.3% with random cropping, whereas our approach experienced a smaller drop of 2.8%. This suggests that our approach learns better semantic features, as it demonstrates greater invariance to natural transformations such as occlusion and variations in object scales. Also, we compare the performance of CLSA (Wang & Qi, 2022) with our approach because we have the same performance after 800 epochs (see Table 1. Note that the CLSA approach uses multi-crop (i.e., five strong and two weak augmentations), while our approach only uses two random crops and the original image. As shown in Table 5 LeOCLR outperforms the CLSA approach by 2.3% after 200 epochs on ImageNet-1K.
Figure 4: Decrease in top-1 accuracy (in % points) of LeOCLR and our own reproduction of Vanilla MoCo-v2 at 200 epochs, under linear evaluation on ImageNet. $R\_Grayscale$ means to remove the grayscale augmentations, and $R\_color$ removes color jitter with grayscale augmentations.
Contrastive instance discrimination approaches are sensitive to the choice of image augmentations (Grill et al., 2020). Thus, we do further analysis of our approach against Moco-v2 (Chen et al., 2020b). These experiments aim to see which model learns better semantic features and creates robust representation under different data augmentations. As shown in Figure 4, both models are affected by removing some data augmentations. However, our approach shows a more invariant representation and less performance impact due to transformation manipulation than vanilla MoCo-v2. For example, when we apply only random cropping augmentation, the performance of vanilla MoCo-v2 is reduced by 28 points (i.e., from 67.5% baseline to 39.5% only random cropping), while our approach reduces only 25 points (i.e., from 71.7% baseline to 46.6% only random cropping). This means our approach learns better semantic features and creates better representation for the given objects than vanilla MoCo-v2.
Refer to caption Refer to caption
(a) Top-1 accuracy (b) Top-5 accuracy
@@
Table 2 presented in Section 4, we fine-tune the representation over the 1% and 10% ImageNet splits from (Chen et al., 2020a) with ResNet-50 architecture. In the ablation study, we compare the fine-tuned representation of our approach and reproduced vanilla MoCo-v2(Chen et al., 2020b) over 1%, 2%, 5%, 10%, 20%, 50%, and 100% of the ImageNet dataset as in (Henaff, 2020; Grill et al., 2020). In this setting, we observed that tuning a LeOCLR representation always outperforms vanilla MoCo-v2. For instance, Figure 5 (a) shows that LeOCLR fine-tuned with 10% of ImageNet labeled data performed better than Vanilla Moco-v2 (Chen et al., 2020b) fine-tuned with 20% of labeled data. This means that our approach is suitable in case we have small labeled data for downstream task than vanilla MoCo-v2.
@@ 6 Conclusion
This paper introduces a new contrastive instance discrimination approach for SSL to enhance representation learning. Our approach alleviates discarding semantic features while attracting two views containing distinct semantic content by incorporating the original image in training. Our method consistently improves the representation learning of contrastive instance discrimination across different benchmark datasets, backbone, and various mechanisms, such as momentum contrast and end-to-end methods. In linear evaluation, we achieved an accuracy of 76.2% on ImageNet after 800 epochs, outperforming several instance discrimination SOTA SSL approaches. Also, we showed the invariance and robustness of our approach through different downstream tasks (i.e., transfer learning and semi-supervised fine-tuning).